OpenStudy (anonymous):

Trying to solve: x^2-11x+30=0

jimthompson5910 (jim_thompson5910):

what pair of factors of 30 add to -11?

OpenStudy (anonymous):

5 &6

OpenStudy (anonymous):

Remember your signs.

OpenStudy (anonymous):

(x+5)(x-6)

jimthompson5910 (jim_thompson5910):

notice 5 times -6 = -30 5 plus -6 = -1

OpenStudy (anonymous):

they need to go the other way around?

OpenStudy (anonymous):

(x+6)(x-5)

OpenStudy (anonymous):

Remember your signs.

OpenStudy (anonymous):

What two values add up to 'negative' eleven. - 11

OpenStudy (anonymous):

-6 + -5

OpenStudy (anonymous):

Correct they are both negative now re-write your previous factoring

OpenStudy (anonymous):

(x-6)(x-5)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Now you can set the 2 different terms equal to 0 separately and solve for x.

OpenStudy (anonymous):

Another method you could use to solve a Quadratic Equation in the form \[ax^2+bx+c = 0\] Is the Quadratic Formula: \[x =\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

you dont have to go that far. lol you can say that x = 6 and 5

OpenStudy (anonymous):

thank you both